DIRC Operations Weekly Report

Period: Wednesday 2007/7/11 - Wednesday 2007/7/18


Oncall: Justine Serrano
Fast Monitoring: Justine Serrano, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~9.7 x 1033 (last week: ~1.02 x 1034)

PEP delivery: 3.19 fb-1 (last weeks: 3.92 fb-1)

BaBar integrated lumi:3.12 fb-1 (last weeks: 3.8 fb-1)

Run 6 total delivery: 61.48 fb-1

Run 6 total BaBar recorded: 59.14 fb-1 total, including 4.18 fb-1 off-peak data

PEP activities/issues this week

  • 07/10: ROD to fix a vacuum leak. After that PEP did scrubbing during several shifts.
  • 07/16: 1 hour and half access to fix a magnet in IR8.
  • 07/17: No abort gap test. Unsuccesful because they didn't manage to have stable beam while filling the gap.

    BaBar activities/issues this week

  • 07/17: several problem with DCH chiller. Karl replaced the filter and the pump.

    DIRC activities/issues this week

  • DIRC backgrounds have been good after the vacuum leak was fixed by PEP.
  • Integrated scaler saturation seen recently was due to a small gate in Jerry's electronics. Jerry is investigating.
  • 07/13: update the white board of the gaz system and update the DQM and pilot instructions.
  • Jochen Kaminski testing his new FEX software.

    See DIRC logbook for details

    Lumi and Background this week

    Scheduled activities

    Pending Work

    Manpower situation

    Justine, Jose and Rolf rotating backup & on-call commissioning duties.

    Links